• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT

    User name R. Snyder

    Log entry time 00:01:09 on July25,2005

    Entry number 149343

    keyword=Swing Shift Summary

    16:00 - 19:50 NO beam. Recovering from CHL issues

    19:50 Attempted HARP scan --didn't work (press start and nothing happened)

    19:55 Raster check (MCC setpoints are 14x14) for approximate 7x7 mm^2 raster

    20:00 Started sieve in septum-scan data collection with these changes

    -All data taken here at 10 uA

    septum current change target

    0% Single C12 foil
    0% 24 cm Optics target
    10% 24 cm Optics target
    10% Single C12 foil
    20% Single C12 foil
    20% 24 cm Optics target

    22:10 Bryan M. makes hall access to rotate the sieves out of the way


    22:35 Lindsay G. moves target to Helium

    22:40 Ask MCC for 0.4 uA, but the beam drifted up to 3 uA and the
    VDC's tripped

    23:10 VDC reboot complete, ask MCC for beam, beam mistakenly drifts
    again to 3 uA and the right VDC's trip off

    23:20 Reboot of right VDC's complete (set to 0V). we ask MCC for
    0.4 uA. After making sure it is stable, we ramp VDC's back up
    4 V.

    23:35 VDC's back to 4, and Bryan M. begins HRS- momentum test
    at different septum currents and different HRS momentum settings

    23:55 set momentum of both spectrometers up 0.002 units.
    Right HRS to 2.738 and Left HRS to 2.734